OpenStudy (anonymous):

Solve the following system of equations by the substitution method. x - y = 0 x - y - 2 = 0 What is the solution set? {(1, -1)} {(-1, 1)} Ø @johnweldon1993

OpenStudy (anonymous):

first, set one equation equal to a variable (solve for x or y). x-y=0 ... solving for x gives you x=y now substitute this into the next equation: x-y-2=0 substituting gives you x-x-2=0. this results in -2=0

OpenStudy (anonymous):

thats what i got but thats not a answer choice @JeanetteBaker

OpenStudy (anonymous):

is "no solution" an option?

OpenStudy (anonymous):

yes , Solve the following system of equations. 2x + y = 3 x = 2y - 1 @JeanetteBaker

OpenStudy (anonymous):

can someone help me??

OpenStudy (anonymous):

@BlueSlurples im trying to get help myself lol

OpenStudy (anonymous):

@brittnicoleee_ ikr lol

OpenStudy (anonymous):

@BlueSlurples maybe i can help you whats the problem

OpenStudy (anonymous):

George drove 12 miles north and 12 mile west. Using a coordinate plane to represent his trip, let the y-axis run north and south, let the x-axis run east and west, and let his original position be the intersection of the axes. Which ordered pair best represents his final location? A.) 12, -12 B.) -12,-12 C.) 12,12 D.)-12,12

OpenStudy (anonymous):

im thinking it A @BlueSlurples

OpenStudy (anonymous):

omfg i love you

OpenStudy (johnweldon1993):

@BlueSlurples yes it is indeed A....you would go up 12......and go left 12...making it 12,-12 @brittnicoleee_ idk if you ever got your answer 2x + y = 3 x = 2y - 1 you know what x equals...so plug that into the first equation 2(2y - 1) + y = 3 4y - 2 + y = 3 5y - 2 = 3 5y = 5 y = 1 if y = 1...plug that into the original first equation 2x + 1 = 3 2x = 2 x = 1 check them both in that second equation x = 2y - 1 1 = 2(1) - 1 1 = 2-1 1 = 1 correct...so x = 1 and y = 1 :)